# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

### v1

@@ 4.2 Existence of time-consistent equilibrium under trading constraint
Under the trading constraint $u_{t}=\sigma_{t}\pi_{t}\in\sigma_{t}A$, where $A\subset\mathbb{R}$ is a bounded closed convex Borel set, assume that (Y,Y~,Z,Z~)($\tilde{Y},$$\tilde{Z}$) is a bounded BMO solution to the BSDE (4.12), then
$\hat{u}_{t}=\sigma_{t}\hat{\pi}_{t}=\mathcal{P}_{\sigma_{t}A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}\right)$ (4.13)
@@
It is clear that $\hat{\pi}\in\Pi_{A}$. We now prove that $\hat{\pi}$ is an equilibrium strategy. Fix $t\in[0,T)$ and let $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;A)$; denote the corresponding perturbed strategy by $\pi^{t,\varepsilon}$. Put $\Delta_{s}=\sigma_{s}(a-\hat{\pi}_{s})$. Then, by (4.9), it holds almost surely that
@@
To verify that the candidate strategy is a time-consistent equilibrium, we need to show that
@@
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right)Z_{s}\right|\mathrm{d}s\right]=0,\quad a.s.,$
Note that $\left\{e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}\right\}_{s\in[t,T]}$ satisfies
@@
Because $\hat{\pi}$ is bounded, by Theorem 3.4.3 in zhang_backward_2017 and the same argument of Lemma 2.7, we deduce that
@@
Moreover, since $Z\in H_{\textup{BMO}}$, $\int_{t}^{T}|Z_{s}|^{2}\mathrm{d}s<\infty,\quad a.s.$ and for fixed $t$ $\int_{t}^{t+\varepsilon}|Z_{s}|^{2}\mathrm{d}s\;\rightarrow\;0\quad a.s..$ By the Conditional Dominated Convergence Theorem
@@
$\leq$ $\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right|^{2}\mathrm{d}s\right]^{\frac{1}{2}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq$ $\frac{1}{\sqrt{\varepsilon}}\mathbb{E}_{t}\left[\sup_{s\in[t,t+\varepsilon]}\left|e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}-1\right|^{2}\right]^{\frac{1}{2}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq$ $C\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}\;\longrightarrow 0,\quad a.s..$
@@
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta Y_{s}}\left(\theta_{s}\Delta_{s}-\frac{1}{2}\Delta_{s}^{2}-(\zeta+1)\sigma_{s}\eta\hat{u}_{s}-\zeta\rho\Delta_{s}Z_{s}\right)\right$.
$\left.-\frac{\gamma}{2}\Delta_{s}^{2}-\gamma\Delta_{s}(\hat{u}_{s}+\rho\tilde{Z}_{s})\mathrm{d}s\right]$
[…]
$\left.\left.+\Delta_{s}\left[e^{-\zeta Y_{s}}\left(\theta_{s}-(\zeta+1)\hat{u}_{s}-\zeta\rho Z_{s}\right)-\gamma(\hat{u}_{s}+\rho\tilde{Z}_{s})\right]\right)\mathrm{d}s\right].$
@@
Thus $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{R}_{t},\pi^{t,\varepsilon,a})-J(t,\hat{R}_{t},\hat{\pi})\right)\leq 0$ , a.s., which completes the proof. ∎
@@ 5 Approximate Time-Consistent Equilibrium for Small $\rho\neq 0$
For $\hat{\pi}\in\Pi_{0}$, $\hat{\pi}$ is called an approximate time-consistent equilibrium strategy with an approximation error $M$ if
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{J(t,\hat{\pi}^{t,\varepsilon,\eta})-J(t,\hat{\pi})}{\varepsilon}\leq M\quad a.s.,$
for any $t\in[0,T)$ and any $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$. Moreover, if $M=O(|\rho|^{\alpha})$, we say that the approximation error [^5] The order of the approximation error effectively characterizes the convergence rate of the equilibrium approximation as $\rho$ tends to 0. is of the order $O(|\rho|^{\alpha})$.
Recall that for $\rho=0$, the equilibrium strategy follows the form of $\hat{u}=\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma}$, where $Y$ is the solution to the BSDE (4.4). For the case $\rho\neq 0$, we consider a trading strategy of the same structure, except that the process $Y$ is replaced by the solution $Y^{\rho}$ to the following BSDE, Specifically, let $Y^{\rho}$ and $\tilde{Y}^{\rho}$ satisfy
@@
Next, we will rigorously verify that the constructed strategy $u^{\rho}$ is an approximate time-consistent equilibrium satisfying Definition 5.1.
Let $R_{t}^{\rho}$ denote the log return at $t$ under strategy $u^{\rho}$. We first show that the Markovian BSDE system (5.1) admits a solution such that $Y^{\rho}$ and $Z^{\rho}$ are uniformly bounded with bounds independent of $\rho\in[-1,1]$ To this end, we need to impose an additional assumption as below.
@@
    There exists a positive constant $\beta>0$ such that $v,v_{x},m,r,\theta$ are Hölder continuous w. r. t $x,t$ with exponents $\beta,\beta/2$ respectively for any bounded cylinder $Q_{T}$ of the layer $R_{T}$.
By the nonlinear Feynman-Kac formula, if the PDE
$\mathcal{L}f=\zeta\rho f_{x}vu^{\rho}+\frac{\zeta}{2}f_{x}^{2}v^{2}-r-\theta u^{\rho}+\frac{\zeta+1}{2}(u^{\rho})^{2}$
admits a classical solution with the terminal conditionf(T,x)=1f(T,x)=1. Let $Y_{t}:=f(t,X_{t})$ and $Z_{t}:=f_{x}(t,X_{t})v(t,X_{t})$. Then $(Y,Z)$ is the solution to our BSDE. This is a cauchy problem and we introduce the following classical result for the quasi-linear equations with the principal part in divergence form,
$u_{t}-\frac{d}{dx_{i}}a_{i}(t,x,u,u_{x})+a(t,x,u,u_{x})=0,$
with $u(0,x)=\psi_{0}(x)$ in $R_{T}=[0,T]\times\mathbb{R}^{n}$ and
$a_{ij}(x,t,u,p)\equiv\frac{\partial a_{i}(x,t,u,p)}{\partial p_{j}},\quad A(x,t,u,p)\equiv a(x,t,u,p)-\frac{\partial a_{i}}{\partial u}p_{i}-\frac{\partial a_{i}}{\partial x_{i}}.$
Remark 5.3.
@@ 1. $\mathrm{a)}$
    $\psi_{0}(x)\in H^{2+\beta}(\Omega)$ in any $\Omega\subset\mathbb{R}^{n}$ and $\max_{\mathbb{R}^{n}}|\psi_{0}(x)|<\infty.$
@@ 2. $\mathrm{b)}$
    $\mu_{1}\xi^{2}\leq a_{ij}(t,x,u,p)\xi^{2}\leq\mu_{2}\xi^{2},$ (5.2)
    with positive constants $\mu_{1}>0$ and $\mu>0$ and
    $A(x,t,u,0)u\geq-b_{1}u^{2}-b_{2},\quad\text{with constants}\ b_{1},\,b_{2}\geq 0.$
@@ 3. $\mathrm{c)}$
    For any bounded cylinder $Q_{T}$ of the layer $R_{T}$ and $|u|\leq M$, where $M$ is a constant depending only on $b_{1},b_{2}$ and $\psi_{0}$, it holds that _i=1^n ( — a_i — + — aiu — ) (1 + —p—) + _i, j=1^n — aixj — + —a— μ(1 + —p—)^2. Moreover, the functions $a_{i}(x,t,u,p)$ and $a(x,t,u,p)$ are continuous, and $a_{i}(t,x,u,p)$ is differentiable w.r.t. variables $x,u,p$ for each $i$.
@@ 4. $\mathrm{d)}$
    For any bounded cylinder $Q_{T}$ of the layer $R_{T}$ and $|u|\leq M$, $|p|\leq M_{1}$, where $M_{1}$ is a constant depending only on $M,\mu,\mu_{1},\mu_{2}$ and $\psi$, the function $a_{i},a,\frac{\partial a_{i}}{\partial p_{j}},\frac{\partial a_{i}}{\partial u}$, and $\frac{\partial a_{i}}{\partial x_{i}}$ are continuous functions satisfying a Hölder condition in $x,t,u$ and $p$ with exponents $\beta,\beta/2,\beta$ and $\beta$ respectively.
Then there exists at least one solution $u(x,t)$ of the Cauchy problem(5.2) in the strip $R_{T}$ that $|u|\leq M$, $|u_{x}|\leq M_{1}$ and belongs to $H^{2+\beta,1+\beta/2}(\overline{Q_{T}})$ for any bounded cylinder $Q_{T}\subset R_{T}$. It will be an element of $H^{2+\beta,1+\beta/2}(R_{T})$ if it is further assumed that the constants in conditions c)-d) do not depend on $Q_{T}$.
@@
Next, we verify that our equations satisfy these conditions under Assumptions 4.1, 4.2 and 5.2. We reverse the time and still denote f(Tt,x)f(T-t,x) by $f(t,x)$ and rewrite the PDE in the divergence form
@@
with $f(0,x)=1$. Here we suppress the dependence of the functions on $(t,x)$ for notational simplicity and also denote that
@@
and the first two conditions are verified given the differentiability of $v$. For condition c), we have
$\left(\left|a_{i}\right|+\left|\frac{\partial a_{i}}{\partial f}\right|\right)(1+|p|)+\left|\frac{\partial a_{i}}{\partial x_{j}}\right|+|a|=\frac{1}{2}v^{2}|p|(1+|p|)+|vv_{x}p|+|a(t,x,f,p)|\leq C(1+|p|^{2}),$
as $\theta,v,v_{x},m,u^{\rho}$ are bounded independent of $\rho$ and $0\leq\rho\leq 1$. We emphasize that $C$ can be chosen independently of $\rho$. To verify the final condition in the theorem, we employ a minor result concerning Hölder continuous functions. That is, if $h_{1}$ and $h_{2}$ are two bounded Hölder continuous functions with $\alpha_{1}$ and $\alpha_{2}$ being their Hölder coefficients respectively, then $h_{1}h_{2}$ is a Hölder continuous function with Hölder coefficient $\min(\alpha_{1},\alpha_{2})$. Then condition d) can be easily verified and there exists a solution $f\in C^{2+\beta,1+\beta/2}(Q_{T})$ for any bounded cylinder $Q_{T}$ of the layer $R_{T}$. Moreover, as $b_{1},b_{2},\mu,\mu_{1},\mu_{2}$ are independent of $\rho$, the constant $M$ and $M_{1}$ also do not depend on $\rho$. Then $Z^{\rho}=mf_{x}$ is bounded independently of $\rho$.
For the same reason, we seek the solution to
$\mathcal{L}g=-r-\theta\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\frac{1}{2}\left(\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}\right)^{2},$
with the terminal condition $g(T,x)=0$. Here $f$ is the solution to equation (5.3). By the same argument, $\tilde{Z}^{\rho}$ is also bounded independently of $\rho$.
Next, we turn to show that $\pi^{\rho}$ is an approximate time-consistent equilibrium. Thanks to the concavity of the utility function, it holds that
$J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\leq e^{R_{t}^{\rho}}\mathbb{E}_{t}\left[U^{\prime}(R_{T}^{\rho})\xi_{T}^{t,\varepsilon}\right]+\frac{\gamma}{2}\left(\textup{Var}(R_{T}^{\rho})-\textup{Var}(R_{T}^{t,\varepsilon})\right)$
Here, we use $\pi^{t,\varepsilon}=\pi^{\rho}+\eta\mathds{1}_{s\in[t,t+\varepsilon)}$ to denote the perturbation strategy and let $\xi^{t,\varepsilon}$ be the resulting difference process for any $\eta\in L^{\infty}_{\mathcal{F}_{t}}(\Omega;\mathbb{R})$ and we shall omit the dependence on $\rho$ and $\eta$ for simplicity.
@@
Under Assumptions 4.1, 4.2, 5.2, the strategy $\pi^{\rho}$ is admissible for any fixed $\rho\in[0,1]$. Moreover, $\pi^{\rho}$ is an approximate time-consistent equilibrium with the approximate error of the order $O(\rho^{2})$ such that for any $t\in[0,T]$ and any fixed $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$,
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\right)\leq C\rho^{2}.$
Here $C$ is a positive constant independent of $t$, $\eta$ and $\rho$.
@@
As $\pi^{\rho}$ is a bounded strategy, it follows from the same arguments for Theorem 4.4 that $\pi^{\rho}$ is an admissible strategy. We next show that $\pi^{\rho}$ is an approximate Nash equilibrium strategy. Fix $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$ and use $\pi^{t,\varepsilon}$ to denote the perturbed strategy. By Lemma 3.4 and the same argument in Theorem 4.11, it holds that, a.s.,
$J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})$
$\leq$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta Y^{\rho}_{s}}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}\eta u^{\rho}_{s}-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\mathrm{d}s\right]-\frac{\gamma}{2}\mathbb{E}\left[\sigma_{s}\eta(2u^{\rho}_{s}+2\rho\tilde{Z}^{\rho}_{s}+\sigma_{s}\eta)\mathrm{d}s\right]$
By the same arguments for Theorem 4.4 and Theorem 4.11, we deduce that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(Y^{\rho}_{s}+R^{\rho}_{s}-R^{\rho}_{t}\right)}-e^{-\zeta Y^{\rho}_{s}}\right)\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}^{2}\pi^{\rho}_{s}\eta-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\right|\mathrm{d}s\right]=0,\quad a.s..$

### v2

@@ 4.2 Existence of time-consistent equilibrium under trading constraint
Assume that (Y,Y~,Z,Z~)($\tilde{Y},$$\tilde{Z}$) is a bounded BMO solution to the BSDE (4.11), then
$\hat{\pi}=\sigma^{-1}\hat{u}=\sigma^{-1}\mathcal{P}_{\sigma A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}\right)$ (4.12)
@@
It is clear that $\hat{\pi}\in\Pi_{A}$. We now prove that $\hat{\pi}$ is an equilibrium strategy. Fix $t\in[0,T)$ and $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;A)$ and let $\Delta_{s}:=\sigma_{s}(a-\hat{\pi}_{s})$. Then, by (4.7) and using the same arguments as in Corollary 3.5 and Lemma 4.4, it holds almost surely that
@@
We first show that
@@
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right)Z_{s}\right|\mathrm{d}s\right]=0,\quad a.s..$
Note that $\{e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}\}_{s\in[t,T]}$ satisfies
@@
Because $\hat{\pi}$ is bounded, by Theorem 3.4.3 in zhang_backward_2017 and using the same argument as in the proof of Lemma 2.6, we deduce that
@@
Moreover, because $Z\in H_{\textup{BMO}}$, we have $\int_{t}^{T}|Z_{s}|^{2}\mathrm{d}s<\infty\,\,a.s.$ and $\int_{t}^{t+\varepsilon}|Z_{s}|^{2}\mathrm{d}s\;\rightarrow\;0\,\,a.s.$ as $\epsilon\to 0$. By the Conditional Dominated Convergence Theorem
@@
$\leq$ $\frac{1}{\varepsilon}\left(\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right|^{2}\mathrm{d}s\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]\right)^{\frac{1}{2}}$
$\leq$ $\frac{1}{\sqrt{\varepsilon}}\left(\mathbb{E}_{t}\left[\sup_{s\in[t,t+\varepsilon]}\left|e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}-1\right|^{2}\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]\right)^{\frac{1}{2}}$
$\leq$ $C\left(\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]\right)^{\frac{1}{2}}\;\longrightarrow 0,\quad a.s..$
@@
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\bigg[\int_{t}^{t+\varepsilon}\bigg(e^{-\zeta Y_{s}}\left(\theta_{s}\Delta_{s}-\frac{1}{2}\Delta_{s}^{2}-(\zeta+1)\Delta_{s}\hat{u}_{s}-\zeta\rho\Delta_{s}Z_{s}\right)$
$\quad\quad\quad\quad\quad\quad-\frac{\gamma}{2}\Delta_{s}^{2}-\gamma\Delta_{s}(\hat{u}_{s}+\rho\tilde{Z}_{s})\bigg)\mathrm{d}s\bigg]$
$=$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\bigg[\int_{t}^{t+\varepsilon}\bigg(-\frac{1}{2}\left(e^{-\zeta Y_{s}}+\gamma\right)(\Delta_{s})^{2}$
$\quad\quad\quad\quad\quad\quad+\Delta_{s}\left[e^{-\zeta Y_{s}}\left(\theta_{s}-(\zeta+1)\hat{u}_{s}-\zeta\rho Z_{s}\right)-\gamma(\hat{u}_{s}+\rho\tilde{Z}_{s})\right]\bigg)\mathrm{d}s\bigg].$
@@
Thus $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{\pi}^{t,\varepsilon})-J(t,\hat{\pi})\right)\leq 0$ a.s., which completes the proof. ∎
@@ 5 Approximate Time-Consistent Equilibrium for Small $\rho\neq 0$
For $\hat{\pi}\in\Pi$, $\hat{\pi}$ is called an approximate time-consistent equilibrium strategy with an approximation error $\Xi$ if
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{J(t,\hat{\pi}^{t,\varepsilon,\eta})-J(t,\hat{\pi})}{\varepsilon}\leq\Xi\quad a.s.$
for any $t\in[0,T)$ and any $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$. Moreover, if $\Xi=O(|\rho|^{\alpha})$, we say that the approximation error [^7] The order of the approximation error effectively characterizes the convergence rate of the equilibrium approximation as $\rho$ tends to 0. is of the order $O(|\rho|^{\alpha})$.
Recall that for $\rho=0$, the equilibrium strategy follows the form of $\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma}$, where $Y$ is the solution to the BSDE (4.3). For the case $\rho\neq 0$, we consider a trading strategy of the same structure, except that the process $Y$ is replaced by the solution $Y^{\rho}$ to the following BSDE, Specifically, let $Y^{\rho}$ and $\tilde{Y}^{\rho}$ satisfy
@@
Next, we will rigorously verify that the constructed strategy $\pi^{\rho}$ is an approximate time-consistent equilibrium satisfying Definition 5.1.
Let $R_{t}^{\rho}$ denote the log return at $t$ under strategy $u^{\rho}$. We first show that the Markovian BSDE system (5.1) admits a solution such that $Y^{\rho}$ and $Z^{\rho}$ are uniformly bounded with bounds independent of $\rho\in[-1,1]$. To this end, we need to impose an additional assumption as below.
@@
    There exists a positive constant $\beta>0$ such that $v,v_{x},m,r,\theta$ are Hölder continuous w. r. t $t,x$ with exponents $\beta/2,\beta$ respectively in $\mathcal{R}_{T}=[0,T]\times\mathbb{R}$.
Let $\varphi(f):=\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}$ for any function $f$ on $\mathcal{R}_{T}$. Suppose that the PDE
$\left\{\begin{aligned} &f_{t}+mf_{x}+\frac{1}{2}v^{2}f_{xx}=\zeta\rho f_{x}v\varphi(f)+\frac{\zeta}{2}f_{x}^{2}v^{2}-r-\theta\varphi(f)+\frac{\zeta+1}{2}\varphi^{2}(f),\\ &f(T,x)=1,\end{aligned}\right$. (5.2)
admits a classical solution $f^{\rho}$. Given such an $f^{\rho}$, consider the linear PDE
$\left\{\begin{aligned} &g_{t}+mg_{x}+\frac{1}{2}v^{2}g_{xx}=-r-\theta\varphi(f^{\rho})+\frac{1}{2}\varphi^{2}(f^{\rho}),\\ &g(T,x)=0,\end{aligned}\right$. (5.3)
and suppose that it admits a classical solution $g^{\rho}$. Then, by setting $Y_{t}^{\rho}:=f^{\rho}(t,X_{t}),\;\tilde{Y}_{t}^{\rho}=g^{\rho}(t,X_{t})$ and $Z_{t}^{\rho}:=f^{\rho}_{x}(t,X_{t})v(t,X_{t}),\;\tilde{Z}^{\rho}=g^{\rho}_{x}(t,X_{t})v(t,X_{t})$, the nonlinear Feynman–Kac formula implies that $(Y^{\rho},\tilde{Y}^{\rho},Z^{\rho},\tilde{Z}^{\rho})$ solves the BSDE (5.1).
Once the solution to (5.2) is obtained, the subsequent PDE (5.3) becomes linear. Therefore, our analysis will primarily focus on the quasi-linear PDE (5.2). To this end, we introduce the following classical existence theorem (Theorem 5.3) for quasi-linear parabolic equations whose principal part is in divergence form. Consider the Cauchy problem
$\left\{\begin{aligned} &u_{t}-\sum_{i=1}^{n}\frac{d}{dx_{i}}m_{i}(t,x,u,u_{x})+a(t,x,u,u_{x})=0,\\ &u(0,x)=\psi_{0}(x).\end{aligned}\right$. (5.4)
To state the theorem, we introduce the notation
$m_{ij}(t,x,u,p)\equiv\frac{\partial m_{i}(t,x,u,p)}{\partial p_{j}},\quad A(t,x,u,p)\equiv a(t,x,u,p)-\frac{\partial m_{i}}{\partial u}p_{i}-\frac{\partial m_{i}}{\partial x_{i}}.$
@@ 1. $\mathrm{a)}$
    $\psi_{0}(x)\in C^{2+\beta}(\mathbb{R}^{n})$.
@@ 2. $\mathrm{b)}$
    $\mu_{1}\xi^{2}\leq m_{ij}(t,x,u,p)\xi^{2}\leq\mu_{2}\xi^{2},$ (5.5)
    with positive constants $\mu_{1}>0$ and $\mu_{2}>0$ and
    $A(t,x,u,0)u\geq-b_{1}u^{2}-b_{2},\quad\text{with constants}\ b_{1},\,b_{2}\geq 0.$
@@ 3. $\mathrm{c)}$
    For any $(t,x)\in\mathcal{R}_{T}=[0,T]\times\mathbb{R}^{n}$ and $|u|\leq M$, where $M$ is a constant depending only on $b_{1},b_{2}$ and $\psi_{0}$, there is a constant $\mu>0$ such that _i=1^n ( — m_i — + — miu — ) (1 + —p—) + _i, j=1^n — mixj — + —a— μ(1 + —p—)^2. Moreover, the functions $m_{i}(t,x,u,p)$ and $a(t,x,u,p)$ are continuous, and $m_{i}(t,x,u,p)$ is differentiable w.r.t. variables $x,u,p$ for each $i$.
@@ 4. $\mathrm{d)}$
    For any $(t,x)\in\mathcal{R}_{T}$ and $|u|\leq M$, $|p|\leq M_{1}$, where $M_{1}$ is a constant depending only on $M,\mu,\mu_{1},\mu_{2}$ and $\psi$, the function $m_{i},a,\frac{\partial m_{i}}{\partial p_{j}},\frac{\partial m_{i}}{\partial u}$, and $\frac{\partial m_{i}}{\partial x_{i}}$ are continuous functions satisfying a Hölder condition in $t,x,u$ and $p$ with exponents $\beta/2,\beta,\beta$ and $\beta$, respectively.
Then there exists at least one solution $u(t,x)$ of the Cauchy problem (LABEL:equ:cauchy) in the strip $\mathcal{R}_{T}$ such that $|u|\leq M$, $|u_{x}|\leq M_{1}$, and $u\in C^{1+\beta/2,2+\beta}(\mathcal{R}_{T})$.
@@
Remark 5.4.
Using Theorem 5.3, we obtain the following result concerning the existence and uniqueness of solutions to (5.2) and (5.3).
Proposition 5.5.
Under Assumptions 4.1, 4.2 and 5.2, there exists a solution $f$ to PDE (5.2) and a solution $g$ to PDE (5.3). Moreover, $f,f_{x},g,g_{x}$ is bounded independent of $\rho$.
Proof.
We focus on the proof of PDE (5.2), as PDE (5.3) can be proved in a similar manner. We verify that (5.2) satisfies conditions (a)–(d) in Theorem 5.3. To begin, we reverse the time by denoting f(Tt,x)f(T-t,x) as $f(t,x)$ and then rewrite the PDE in divergence form.
@@
with $f(0,x)=1$. We also denote that
@@
and the first two conditions are verified. For Condition c), we have
$\left(\left|m_{1}\right|+\left|\frac{\partial m_{1}}{\partial f}\right|\right)(1+|p|)+\left|\frac{\partial m_{1}}{\partial x_{j}}\right|+|a|=\frac{1}{2}v^{2}|p|(1+|p|)+|vv_{x}p|+|a(t,x,f,p)|\leq C(1+|p|^{2}),$
as $\theta,v,v_{x},m,\varphi(f)$ are bounded independent of $\rho$ and $1\leq\rho\leq 1$. We emphasize that $C$ can be chosen independently of $\rho$.
To verify the final condition in the theorem, we employ a minor result concerning Hölder continuous functions. That is, if $h_{1}$ and $h_{2}$ are two bounded Hölder continuous functions with $\alpha_{1}$ and $\alpha_{2}$ being their Hölder coefficients respectively, then $h_{1}h_{2}$ is a Hölder continuous function with Hölder coefficient $\min(\alpha_{1},\alpha_{2})$. Then Condition d) can be easily verified and thus there exists a solution $f\in C^{1+\beta/2,\,2+\beta}(\mathcal{R}_{T})$. Moreover, as $b_{1},b_{2},\mu,\mu_{1},\mu_{2}$ are independent of $\rho$, the constants $M$ and $M_{1}$ in Theorem 5.3 are also independent of $\rho$. ∎
By Proposition 5.5 and the nonlinear Feynman–Kac formula, there exists a bounded solution $(Y^{\rho},\tilde{Y}^{\rho},Z^{\rho},\tilde{Z}^{\rho})$ to the BSDE (5.1) and the processes $Z^{\rho}$ and $\tilde{Z}^{\rho}$ are bounded by a constant that does not depend on $\rho$.
Next, we turn to show that $\pi^{\rho}$ is an approximate time-consistent equilibrium.
@@
Under Assumptions 4.1, 4.2, 5.2, $\pi^{\rho}\in\Pi$ for any fixed $\rho\in[-1,1]$. Moreover, $\pi^{\rho}$ is an approximate time-consistent equilibrium with the approximate error of the order $O(\rho^{2})$ such that for any $t\in[0,T]$ and any fixed $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$ [^8] Recall the notations in Remark 2.10 and we shall also omit the dependence of the perturbed strategy on $\rho$ for notational simplicity.,
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\pi^{t,\varepsilon})-J(t,\pi^{\rho})\right)\leq C\rho^{2}.$
Moreover, $C$ is a positive constant independent of $t$, $\eta$ and $\rho$.
@@
As $\pi^{\rho}$ is bounded, it follows from the same arguments used in the proof of Theorem 4.4 that $\pi^{\rho}\in\Pi$. We next show that $\pi^{\rho}$ is an approximate Nash equilibrium strategy. Fix $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$. By Corollary 3.5 and the same line of reasoning as in Theorem 4.4, we obtain, almost surely, that
$J(t,\pi^{t,\varepsilon})-J(t,\pi^{\rho})\leq$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta Y^{\rho}_{s}}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}\eta u^{\rho}_{s}-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\mathrm{d}s\right]$
$-\frac{\gamma}{2}\mathbb{E}_{t}\left[\sigma_{s}\eta(2u^{\rho}_{s}+2\rho\tilde{Z}^{\rho}_{s}+\sigma_{s}\eta)\mathrm{d}s\right].$
Using the same arguments for Theorem 4.4 and Theorem 4.11, we deduce that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(Y^{\rho}_{s}+R^{\rho}_{s}-R^{\rho}_{t}\right)}-e^{-\zeta Y^{\rho}_{s}}\right)\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}^{2}\pi^{\rho}_{s}\eta-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\right|\mathrm{d}s\right]=0\quad a.s..$